I'm not sure if this goes into the finance section, but here is the problem. Adam buys oranges at the rate of three oranges for $1. He sells them at a rate of five oranges for $2. In order to make a profit of at least $10, what is the minimum number of oranges he must sell?
I tried doing proportions but I don't think I did them right.

Answer by rfer(16322)

1/3=0.333
2/5=0.4
0.4-0.333=0.0.67
10/0.67=149.23=150 min